Homework Ch. #2 Problem 2.60 Constants | Periodic Table Part A A stone is thrown vertically upward with a speed of 23.0 m/s How fast is it moving when it reaches a height of 8.10 m? Express your answer in meters per second. UE m/s Submit Part B How much time is required to reach this height? Express your answer in seconds. If there is more than one answer, separate them by a comma t- Submit An
Answer #1

vi = 23 m/s

and a = - 9.8 m/s^2


(A) Using vf^2 - vi^2 = 2 a d

v^2 - (23)^2 = 2(-9.8)(8.10)

v = 19.2 m/s

(B) vf = vi + a t

19.2 = 23 - 9.8t

t = 0.38 sec